OpenStudy (anonymous):

Use combinations of Properties 1 and 2 for radicals to simplify the following problem as much as possible. Assume the variable represents a positive number.

OpenStudy (anonymous):

\[\frac{ 4 \sqrt{18a^2b^2}}{ \sqrt{9} }\]

OpenStudy (anonymous):

@kx2bay

OpenStudy (anonymous):

hey cookie, how are you ? could you move the denominator under the sq.root of the numerator?

OpenStudy (anonymous):

Hi :) I'm good, you? What do you mean?

OpenStudy (anonymous):

\[4\sqrt{(18a ^{2}b ^{2})/9}\]

OpenStudy (anonymous):

then divide 18 by 9, you get 2

OpenStudy (anonymous):

a^2 & b^2 becomes a & b

OpenStudy (anonymous):

then take a^2 b^2 out of the radical and you'll be left with sq.root of 2

OpenStudy (anonymous):

yes that's correct

OpenStudy (anonymous):

What d o I do with the 4?

OpenStudy (anonymous):

it stays, the answer looks like this \[4ab \sqrt{2}\]

OpenStudy (anonymous):

Got it !

OpenStudy (anonymous):

am just not sure what your book says about property 1 and 2 of radicals, need to paste that if the answer isn't what they expect

OpenStudy (anonymous):

I don't even use my book lol it's all online, including the ebook.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

I guess then refer to your online ebook for the properties as I don't have a reference to it no worries cheers

OpenStudy (anonymous):

You want me to look for them & paste them on here?

OpenStudy (anonymous):

ok that will be good, thanks

OpenStudy (anonymous):

found them!!

OpenStudy (anonymous):

what do they state?

OpenStudy (anonymous):

cool, so can you see how we applied both prop's ?

OpenStudy (anonymous):

prop 1 is when you took a^2 and b^2 out of the sq.root and prop.1 when we moved the 9 under the same sq.root with the 19.

OpenStudy (anonymous):

prop 1 is when you took a^2 and b^2 out of the sq.root and prop.1 when we moved the 9 under the same sq.root with the 19.

OpenStudy (anonymous):

Oh, I see it